# Fast-Locking and High-Resolution Mixed-Mode DLL with Binary Search and Dead Clock Detection for Wide Frequency Ranges in 3-nm FinFET CMOS

### v1

@@ Abstract
This paper presents a mixed-mode delay-locked loop (MM-DLL) with binary search (BS) locking, designed to cover a broad frequency range from 533 MHz to 4.2 GHz. The BS locking scheme optimizes the locking time, reducing it from a linear to a logarithmic function, completing in $\log_{2}(2^{B})+1$ cycles, where $B$ represents the digital-to-analog (DAC) resolution controlling the voltage-controlled delay line (VCDL). At the start of the BS process, large step sizes can cause significant bias overshoots, potentially leading to dead clock conditions (i.e., clocks fail to propagate through the VCDL). To address this issue, a toggle detector is introduced to monitor clock activity and adjust the binary search controller. Upon detecting a dead clock, the controller reverts the DAC code to the previous working code and resumes the BS with a reduced step size. Fabricated in a 3-nm FinFET CMOS process, the proposed MM-DLL achieves a locking time of under 10.5 ns while consuming 5.4 mW from a 0.75 V supply at 4.2 GHz. The measured performance includes a high resolution of 0.73 ps, with a static phase error of 0.73 ps, RMS jitter of 1.2 ps, and peak-to-peak jitter of 4.9 ps.
@@ I Introduction
High-speed parallel wireline links are fundamental in any modern system-on-chip (SoC), and at the heart of these interfaces stands the delay-locked loop (DLL). This circuit is specialized in generating one or multiple phases of a reference clock for the deskew systems, and clock and data recovery [1, 2]. With the increasing data rates of double data rate (DDR) standards and recent advancements in die-to-die (D2D) links for 2.5D and 3D integration [3, 4, 5], ultra-wide-range DLLs are required to meet the specifications of new standards while maintaining backward compatibility. Additionally, to achieve day-long battery life in state-of-the-art microprocessors, deep low-power modes are required, making low exit times critical for the overall performance of the SoC.
@@
In recent years, all-digital DLLs (AD-DLL) have been extensively studied due to their superior scalability with CMOS process, lower power consumption [6, 7], and faster locking time [8, 9, 10, 11]. However, AD-DLLs struggle to provide the fine resolution required in high-performance wireline links to minimize static phase errors and jitter [12]. Conversely, analog and mixed-mode DLLs (A-DLLs and MM-DLLs, respectively) can achieve higher operating frequencies and sub-picosecond (ps) resolution by means of highly optimized voltage-controlled delay lines (VCDLs) and phase detectors (PDs) [13, 9]. This sub-ps time resolution is enabled by fine-grained control of bias or power supply of the VCDL, typically generated by high-resolution digital-to-analog converters (DACs) or charge-pumps (CPs). However, the enhanced temporal resolution in A-DLLs and MM-DLLs comes at the cost of increased locking times since the locking duration is inversely proportional to the time resolution [9]. Furthermore, to prevent harmonic locking, i.e., locking at an incorrect multiple of the desired delay between the feedback clock (fbclk) and the reference clock (refclk), the initial condition of the VCDL bias must ensure that fbclk leads refclk at the input of the PD. As a result, the DAC or CP often must be initialized at the minimum code, leading to longer locking times.
@@
A potential solution to long locking times is the use of coarse-fine modes, which adjust the DAC code or capacitive ratios in coarse steps during DLL initialization, until the first polarity change in the phase difference ($\Delta\phi$) between fbclk and refclk is detected by the PD [8]. Once this occurs, the step size is reduced to achieve fine-grained resolution until the locking bias condition is met ($\Delta\phi\approx 0$). While the coarse-fine scheme improves locking times, it can lead to a condition known as “dead clocks,” where the clocks stop propagating through the VCDL due to large bias overshoots. This condition is irreversible as the PD loses track of the polarity of $\Delta\phi$, necessitating to restart the DLL initialization process with adjusted parameters, such as the coarse step size, or the capacitive load of the delay elements (DEs) within the VCDL. Importantly, in many DLL designs, this issue may create a false locking condition and go undetected until later stages of the initialization flow, potentially affecting subsystems that depend on the output clocks of the DLL.
@@
In this work, a fast-locking, wide frequency range, and high-resolution MM-DLL architecture is presented. The proposed architecture, implemented in a 3-nm FinFET CMOS process, achieves a locking time of eleven cycles over a wide frequency range from 533 MHz to 4.2 GHz by means of a binary search (BS) locking scheme. This DLL reduces the locking time by 5-10$\times$ compared to traditional coarse-fine schemes, widely used in MM-DLLs. The BS is implemented by a finite-state machine (FSM) with bounded number of steps; and thus, the locking time becomes deterministic. A deterministic locking time reduces latency during transitions from idle to active modes and minimizes both startup and low-power mode exit times.
To address bias overshoot and potential dead clock events, we introduce a toggle detection circuit that can identify dead clock conditions. The DLL loop controller samples the output of the toggle detector after each code change and applies corrective actions without needing to restart the entire initialization process. Leveraging the excellent time resolution in advanced CMOS process nodes, we achieve a sub-ps resolution and static phase error of 0.73 ps. The presented DLL consumes 5.4 mW from a 0.75-V power supply at 4.2 GHz.
@@ II-A MM-DLL Basic Architecture
A typical MM-DLL is composed by a differential VCDL, a PD, a digital loop control or filter (DLC), a clock divider (÷$\div N$), and a DAC as depicted in Fig. 1. The delay is dynamically controlled by the VCDL to guarantee that the $\Delta\phi$ is locked regardless of PVT and frequency variations [17]. The VCDL generates multiple equally-spaced phases of the input differential clocks (clkin, clkinb). VCDLs are implemented as a chain of matched DEs with varying load capacitance and/or varying driving strength, and generates multiple equally-spaced clock phases (eight in this example, clkout<7:0>). The driving strength is modulated either by changing the power supply or by using a current-starved topology with bias voltage [18]. The DAC generates the bias or supply voltage for the DE ($V_{ctrl}$).
The phase difference $\Delta\phi$ is measured by the PD, which generates a binary output (pder). As an example, pder will be logic ‘1’ if refclk leads and logic ‘0’ otherwise. The PD controls the DLC, which adjusts the DAC code to ensure that $\Delta\phi\approx 0$. Due to the binary nature of pder and the presence of jitter, the DLL will dither around by at least ±$\pm 1$ code even after locking. The DLC is controlled by clkctrl, a subdivided version of the input clock (clkin). Running the DLC at a lower frequency is necessary as synthesizing digital logic at very high frequencies is typically challenging [9], and due to the relatively large settling time of the DAC compared to the input frequency. Consequently, the update rate, or loop bandwidth is primarily determined by the DLC and the DAC.
@@
where $K_{VCDL}$ is the gain of the VCDL ($K_{VCDL}=\partial\Delta\phi/\partial V_{ctrl}$), and $\Delta V_{LSB}$ is the least significant bit (LSB) resolution of the DAC. The voltage resolution of the DAC is $\Delta V_{LSB}=V_{FS}/2^{N}$, where $V_{FS}$ and $N$ are the full-scale voltage and number of bits, respectively. Then, the phase resolution ($\Delta\phi_{LSB}$) is expressed as
@@
where $\Delta\phi_{init}$ is the initial phase difference between the fbclk and the refclk, and $f_{clkctrl}$ is the frequency of clkctrl, respectively. Hence, as expected, $T_{Lock}$ is inversely proportional to $\Delta T_{LSB}$, $f_{clkin}$, and $f_{clkctrl}$, and proportional to $\Delta\phi_{init}$. In general, $f_{clkctrl}$ will be selected to be 4-10$\times$ lower than $f_{clkin}$ [19].
@@ III-A Architecture
Figure 3: (a) Schematic diagram implementation of the delay element (DE) within the VCDL. (b) Current-starved (CS) inverter schematic diagram.
@@
The proposed MM-DLL architecture is illustrated in Fig. 2(a). It comprises a VCDL, bang-bang PD (BBPD), clock divider, DLC, and DAC. To improve locking time, the DLC employs a BS control mechanism, where an FSM generates DAC codes by navigating a binary tree. The BS control includes an accumulator, a shift register to adjust the step size, and registers to store the current and previous DAC codes. The FSM is clocked by a configurable clock divider, which supports divisions of $N=1,2,4,6,$ and $8$. This clock divider reduces the loop bandwidth, enhancing the timing margins of the synchronous circuits within the FSM, and minimizing output clock dithering. A toggle detector samples refclk and raises a flag (toggling) in the occurrence of a dead clock condition.
@@
A 10-bit DAC produces the analog bias signal for the PMOS devices (vctrlp) in the VCDL, while a replica circuit generates the bias for the NMOS devices (vctrln). These two bias voltages modulate the delay of the VCDL. The VCDL generates eight clock outputs (clkout<7:0>), evenly spaced by $T_{clkin}/8$, where $T_{clkin}$ is the period of the input clock. The VCDL is implemented as a chain of ten DEs, as shown in Fig. 2(b). The ninth DE generates fbclk, while the tenth DE acts as a dummy load for delay matching.
@@
The delay is controlled by adjusting the effective resistance and capacitance at each DE stage. Each DE consists of a current-starved (CS) inverter with variable load capacitance, as shown in Fig. 3(a). Cross-coupled inverters are used to maintain the pseudo-differential clock paths aligned, minimizing skew within the VCDL. The load capacitance is regulated by a bank of capacitors, while the effective resistance of the pull-up and pull-down of the CS inverter is modulated by the bias voltages of the PMOS and NMOS tails, as shown in Fig. 2(d) to maintain a balanced duty cycle. Furthermore, the effective resistance of the CS tails can be modulated by enabling or disabling parallel branches by means of nbw[3:0] and pbw[3:0], as shown in Fig. 3(b). This flexible delay stage implementation, coupled with high DAC resolution, enables the DLL to cover a wide frequency range—from 533 MHz to 4.2 GHz—while achieving a $\Delta T_{LSB}$ of less than 1 ps.
III-B Binary Search Locking Procedure
@@
Figure 4: Locking algorithms example. (a) Coarse-fine scheme.
@@
Examples of locking search procedures for the coarse-fine and the BS are depicted in Fig. 4(a) and Fig. 4(b), respectively. The coarse-fine algorithm begins with coarse step sizes of $k=50$ until a polarity change in pder. Next, the search proceeds with step sizes of one until the locking condition is met. In contrast, the BS scheme navigates a binary tree. The step size starts at half the full range (i.e., $1024/2=512$). The step size is halved at each cycle of $f_{clkctrl}$. Once the step size reaches one, the DLL locks. In the absence of dead clock conditions, the locking process is completed in $\log_{2}{\left(2^{B}\right)}+1=B+1$ steps, where $B$ is the bit precision of the DAC ($B=10$ in this case). Therefore, the locking time becomes $T_{Lock}=(B+1)/f_{clkctrl}$, which is 5-10$\times$ faster than the coarse-fine scheme.
@@
For comparison, the proposed MM-DLL is tested with both the coarse-fine scheme and the BS algorithm. The simulation results of the startup locking process using the coarse-fine algorithm are shown in Fig. 5(a) and Fig. 5(b) for input frequencies of 0.8 GHz and 4.2 GHz, respectively, across three process corners. In particular, the skew between refclk and fbclk ($\Delta T$) initially rises sharply due to the coarse steps, resulting in significant overshoot. When the polarity of pder changes, finer steps are introduced until the DLL reaches lock, with $\Delta T$ dithering around zero in steady state. This dithering is the result of the binary output of the BBPD, which causes the DAC code to fluctuate by at least $\pm$1 code even in the locked state. Note the long $T_{Lock}$ in both cases, particularly at 800 MHz, the locking process can exceed 140 ns.
@@
Simulations of the BS scheme for three corners are depicted in Fig. 6(a) and Fig. 6(b) for 0.8 GHz and 4.2 GHz, respectively. It can be observed that the DLL achieves lock within approximately 12.5 ns and 10.5 ns, for 0.8 GHz and 4.2 GHz, respectively, which corresponds to eleven cycles of clkctrl (the simulation includes an initial reset and enable). This represents a reduction of 5-10x in the locking time compared to the coarse-fine scheme. Notably, the startup time remains consistent across PVT conditions, as the BS ensures deterministic locking time in the absence of dead clock events. The deterministic locking time is a key advantage of the proposed architecture. Even in worst-case scenarios with dead clock conditions, the maximum locking time remains predictable. Additionally, the reduced locking time significantly enhances the cold boot and low-power mode exit latency in microprocessors, as well as in high-performance parallel interfaces.
@@
Figure 5: Simulation of the coarse-fine locking scheme for different PVT conditions. Change in $\Delta T$ for $f_{clkin}$ of (a) 800 MHz and (b) 4.2 GHz, at nominal (blue), fast (red), and slow (black) corners.
@@
Figure 6: Simulation results of the BS locking scheme for different PVT conditions. Change of $\Delta T$ for $f_{clkin}$ of (a) 0.8 GHz and (b) 4.2 GHz, at nominal (blue), fast (red), and slow (black) operating conditions.
III-C Dead Clock or Toggle Detector

### v2

# Fast-Locking and High-Resolution Mixed-Mode DLL with Binary Search and Clock Failure Detection for Wide Frequency Ranges in 3-nm FinFET CMOS

@@ Abstract
[…]
@@ I Introduction
High-speed parallel wireline links are fundamental in any modern system-on-chip (SoC), and at the heart of these interfaces stands the delay-locked loop (DLL). This circuit is specialized in generating one or multiple phases of a reference clock for the deskew systems, and clock and data recovery [1, 2]. With the increasing data rates of double data rate (DDR) standards and recent advancements in die-to-die (D2D) links for 2.5D and 3D integration [3, 4, 5, 6], ultra-wide-range and low-latency DLLs are required to meet the specifications of new standards while maintaining backward compatibility. Additionally, to achieve day-long battery life in state-of-the-art microprocessors, deep low-power modes are required, making low exit times critical for the overall performance of the SoC.
@@
In recent years, all-digital DLLs (AD-DLL) have been extensively studied due to their superior scalability with CMOS process, lower power consumption [7, 8], and faster locking time [9, 10, 11, 12]. However, AD-DLLs struggle to provide the fine resolution required in high-performance wireline links to minimize static phase errors and jitter [13]. Conversely, analog and mixed-mode DLLs (A-DLLs and MM-DLLs, respectively) can achieve higher operating frequencies and sub-picosecond (ps) resolution by means of highly optimized voltage-controlled delay lines (VCDLs) and phase detectors (PDs) [14, 10]. This sub-ps time resolution is enabled by fine-grained control of bias or power supply of the VCDL, typically generated by high-resolution digital-to-analog converters (DACs) or charge-pumps (CPs). However, the enhanced temporal resolution in A-DLLs and MM-DLLs comes at the cost of increased locking times since the locking duration is inversely proportional to the time resolution [10]. Furthermore, to prevent harmonic locking, i.e., locking at an incorrect multiple of the desired delay between the feedback clock (CLKFB) and the reference clock (CLKREF), the initial condition of the VCDL bias must ensure that CLKFB leads CLKREF at the input of the PD. As a result, the DAC or CP often must be initialized at the minimum code, leading to longer locking times.
@@
A potential solution to long locking times is the use of coarse-fine modes, which adjust the DAC code or capacitive ratios in coarse steps during DLL initialization, until the first polarity change in the phase difference ($\Delta\phi$) between CLKFB and CLKREF is detected by the PD [9]. Once this occurs, the step size is reduced to achieve fine-grained resolution until the locking bias condition is met ($\Delta\phi\approx 0$). While the coarse-fine scheme improves locking times, it can lead to a clock failure condition, where the clocks stop propagating through the VCDL due to large bias overshoots. This condition is irreversible as the PD loses track of the polarity of $\Delta\phi$, necessitating to restart the DLL initialization process with adjusted parameters, such as the coarse step size, or the capacitive load of the delay elements (DEs) within the VCDL. Importantly, in many DLL designs, this issue may create a false locking condition and go undetected until later stages of the initialization flow, potentially affecting subsystems that depend on the output clocks of the DLL.
@@
In this work, a fast-locking, wide frequency range, and high-resolution MM-DLL architecture is presented. The proposed architecture, implemented in a 3-nm FinFET CMOS process, achieves a locking time of eleven cycles over a wide frequency range from 533 MHz to 4.26 GHz by means of a binary search (BS) locking scheme. This DLL reduces the locking time by 5-10$\times$ compared to traditional coarse-fine schemes, widely used in MM-DLLs. The BS is implemented by a finite-state machine (FSM) with bounded number of steps; and thus, the locking time becomes deterministic. A deterministic locking time reduces latency during transitions from idle to active modes and minimizes both startup and low-power mode exit times.
To address bias overshoot and potential clock failure events, we introduce a toggle detection circuit that can identify stalled clock conditions. The DLL loop controller samples the output of the toggle detector after each code change and applies corrective actions without needing to restart the entire initialization process. Leveraging the excellent time resolution in advanced CMOS process nodes, we achieve a sub-ps resolution and static phase error of 0.73 ps. The presented DLL consumes 5.4 mW from a 0.75-V power supply at 4.26 GHz.
@@ II-A MM-DLL Basic Architecture
A typical MM-DLL is composed by a differential VCDL, a PD, a digital loop control or filter (DLC), a clock divider (÷$\div N$), and a DAC as depicted in Fig. 1. The delay is dynamically controlled by the VCDL to guarantee that the $\Delta\phi$ is locked regardless of PVT and frequency variations [18]. The VCDL generates multiple equally-spaced phases of the input differential clocks (CLKIN,P, CLKIN,N). VCDLs are implemented as a chain of matched DEs with varying load capacitance and/or varying driving strength, and generates multiple equally-spaced clock phases (eight in this example, CLKOUT[7:0]). The driving strength is modulated either by changing the power supply or by using a current-starved topology with bias voltage [19]. The DAC generates the bias or supply voltage for the DE (VCTRL).
The phase difference $\Delta\phi$ is measured by the PD, which generates a binary output (PDER). As an example, PDER will be logic ‘1’ if CLKREF leads and logic ‘0’ otherwise. The PD controls the DLC, which adjusts the DAC code to ensure that $\Delta\phi\approx 0$. Due to the binary nature of PDER and the presence of jitter, the DLL will dither around by at least ±$\pm 1$ code even after locking. The DLC is controlled by CLKCTRL, a subdivided version of the input clock (CLKIN). Running the DLC at a lower frequency is necessary as synthesizing digital logic at very high frequencies is typically challenging [10], and due to the relatively large settling time of the DAC compared to the input frequency. Consequently, the update rate, or loop bandwidth is primarily determined by the DLC and the DAC.
@@
where $K_{VCDL}$ is the gain of the VCDL ($K_{VCDL}=\partial\Delta\phi/\partial V_{ctrl}$), and $\Delta V_{LSB}$ is the least significant bit (LSB) resolution of the DAC. The DAC voltage resolution is $\Delta V_{LSB}=V_{FS}/2^{N}$, where $V_{FS}$ and $N$ are the full-scale voltage and number of bits, respectively. Then, the phase resolution ($\Delta\phi_{LSB}$) is expressed as
@@
where $\Delta\phi_{init}$ is the initial phase difference between the CLKFB and the CLKREF, and $f_{clkctrl}$ is the frequency of CLKCTRL, respectively. Hence, as expected, $T_{Lock}$ is inversely proportional to $\Delta T_{LSB}$, $f_{clkin}$, and $f_{clkctrl}$, and proportional to $\Delta\phi_{init}$. In general, $f_{clkctrl}$ will be selected to be 4-10$\times$ lower than $f_{clkin}$ [20].
@@ III-A Architecture
Figure 3: Schematic diagram of (a) the pseudo-differential delay element (DE) within the VCDL, and (b) the pseudo-differential current-starved inverter (CSI).
@@
The proposed MM-DLL architecture is illustrated in Fig. 2(a). It comprises a differential VCDL, bang-bang PD (BBPD), clock divider, DLC, and DAC. To improve locking time, the DLC employs a BS control mechanism, where an FSM generates DAC codes by navigating a binary tree. The BS control includes an accumulator, a shift register to adjust the step size, and registers to store the current and previous DAC codes. The FSM is clocked by a configurable clock divider, which supports divisions of $N=1,2,4,6,$ and $8$. This clock divider reduces the loop bandwidth, enhancing the timing margins of the synchronous circuits within the FSM, and minimizing output clock dithering. A toggle detector samples CLKREF and raises a flag (toggling) in the occurrence of a stalled clock condition.
@@
A 10-bit DAC produces the analog bias signal for the PMOS devices (VCTRLP) in the VCDL, while a replica circuit generates the bias for the NMOS devices (VCTRLN). These two bias voltages modulate the delay of the VCDL. The VCDL generates eight clock outputs (CLKOUT[7:0]), evenly spaced by $T_{clkin}/8$, where $T_{clkin}$ is the period of the input clock. The VCDL is implemented as a chain of ten DEs, as shown in Fig. 2(b). The ninth DE generates CLKFB, while the tenth DE acts as a dummy load for delay matching. The number of DEs is kept to the possible minimum, as the output jitter in DLLs is directly proportional to the number of DEs in the VCDL [21].
III-B Delay Element Design
The DE consists of two of cascaded pseudo-differential current-starved inverters (PS-CSIs), each loaded by a bank of capacitors, as illustrated in Fig. 3(a). The total load capacitance ($C_{B}$) is adjusted by a bank of capacitors, controlled by CBEN[1:0]. The output of the second PS-CSI stage drives a tri-state buffer, enabled by ENCLKOUT,P and ENCLKOUT,N, for the CLKOUT,P and CLKOUT,N, respectively. The DE outputs (DEOUT,P and DEOUT,N) drive the next DE stage. The first PS-CSI stage is loaded by dummy tri-state buffers for delay matching. Since both PS-CSIs are matched and evenly loaded, the total delay the DE, $t_{d,DE}$, is the sum of the delays of both CSIs, i.e., $t_{d,DE}=2t_{d,CSI}$, where $t_{d,CSI}$ is the delay of the CSI.
Each PS-CSI consists of two main inverters, each having two current-limiting transistors (i.e., tails), NMOS-based and PMOS-based, which starve the current of the main inverters, regulating their drive strength. The tails are controlled by bias voltages VCTRLP and VCTRLN, as shown in Fig. 3(b). These biases are nearly symmetrical to ensure a balanced duty cycle. Thus, $t_{d,CSI}$ depends on the effective resistance ($R_{eff}$) and $C_{B}$, given by:
$t_{d,CSI}\propto R_{eff}C_{B}~{},$ (5)
where $R_{eff}$ is modeled as:
$R_{eff}=\frac{L/W}{k(V_{DD}-V_{T})}+\frac{L/W_{tail}}{k(V_{CTRLN}-V_{T})}.$ (6)
Here, $L$, $W$, $k$, and $V_{T}$ represent the transistor length, width, transconductance, and threshold voltage, respectively, while $W_{tail}$ denotes the width of the tail transistors.
The number of active tail branches is tunable via BWN[3:0] and BWP[3:0], which selectively connect each branch to $V_{DD}$ ($V_{SS}$) or to VCTRLP (VCTRLN) for the pull-up (down) paths [see Fig. 3(b)]. A static branch (not shown) is always active to provide a minimum pull-up/down path for the main inverters. Thus, the values of BWN[3:0] and BWP[3:0] tune $W_{tail}$. Their values are dynamically adjusted according to the PVT and $f_{clkin}$ operating conditions.
@@ III-B Delay Element Design
Cross-coupled inverters are used to maintain the pseudo-differential clock paths aligned, minimizing skew within the VCDL. These cross-coupled inverters are chosen to be six times smaller than the main inverters. This flexible delay stage implementation, coupled with high DAC resolution, enables the DLL to cover a wide frequency range—from 533 MHz to 4.26 GHz—while achieving a $\Delta T_{LSB}$ of less than 1 ps.
III-C Binary Search Locking Procedure
@@
Figure 4: Locking algorithms example. (a) Coarse-fine scheme. (b) BS scheme.
@@
Examples of locking search procedures for the coarse-fine and the BS are depicted in Fig. 4(a) and Fig. 4(b), respectively. The coarse-fine algorithm begins with coarse step sizes of $k=50$ until a polarity change in PDER. Next, the search proceeds with step sizes of one until the locking condition is met. In contrast, the BS scheme navigates a binary tree. The step size starts at half the full range (i.e., $1024/2=512$). The step size is halved at each cycle of $f_{clkctrl}$. Once the step size reaches one, the DLL locks. In the absence of clock failure conditions, the locking process is completed in $\log_{2}{\left(2^{B}\right)}+1=B+1$ steps, where $B$ is the bit precision of the DAC ($B=10$ in this case). Therefore, the locking time becomes $T_{Lock}=(B+1)/f_{clkctrl}$, which is 5-10$\times$ faster than the coarse-fine scheme.
@@
For comparison, the proposed MM-DLL is tested with both the coarse-fine scheme and the BS algorithm. The simulation results of the startup locking process using the coarse-fine algorithm are shown in Fig. 5(a) and Fig. 5(b) for input frequencies of 800 MHz and 4.26 GHz, respectively, across three process corners. In particular, the skew between CLKREF and CLKFB ($\Delta T=2\pi f_{clkin}\Delta\phi$) initially rises sharply due to coarse steps, resulting in significant overshoot. When the polarity of PDER changes, finer steps are introduced until the DLL reaches lock, with $\Delta T$ dithering around zero in steady state. This dithering arises from the binary nature of the BBPD output, causing the DAC code to fluctuate by at least $\pm$1 code even in the locked state. Additionally, the $T_{Lock}$ is relatively long at both frequencies. Notably at 800 MHz, the locking process can extend beyond 140 ns, making it impractical for latency-critic applications.
@@
Simulations of the BS scheme for three corners are depicted in Fig. 6(a) and Fig. 6(b) for 800 MHz and 4.26 GHz, respectively. It can be observed that the DLL achieves lock within approximately 12.5 ns and 10.5 ns, for 800 MHz and 4.26 GHz, respectively, which corresponds to eleven cycles of CLKCTRL (the simulation includes an initial reset and enable). This represents a reduction of 5-10$\times$ in the locking time compared to the coarse-fine scheme. Notably, the startup time remains consistent across PVT conditions, as the BS ensures deterministic locking time in the absence of clock failure events. The deterministic locking time is a key advantage of the proposed architecture. Even in worst-case scenarios with stalled clock conditions, the maximum locking time remains predictable. Additionally, the reduced locking time significantly enhances the cold boot and low-power mode exit latency in microprocessors, as well as in high-performance parallel interfaces, such as D2D, where latency is a critical design consideration.
@@
Figure 5: Simulation of the coarse-fine locking scheme for different PVT conditions. Change in $\Delta T$ for $f_{clkin}$ of (a) 800 MHz and (b) 4.26 GHz, at nominal (blue), fast (red), and slow (black) corners.
@@
Figure 6: Simulation results of the BS locking scheme for different PVT conditions. Change of $\Delta T$ for $f_{clkin}$ of (a) 800 MHz and (b) 4.26 GHz, at nominal (blue), fast (red), and slow (black) operating conditions.
III-D Clock Failure Detector